OpenStudy (anonymous):

Find tan x if cos x = 4/9 and 0 < x < pi/2 (posting in equation editor)

OpenStudy (anonymous):

\[\tan x = \frac{ \sin x }{ \cos x }\] \[\cos x = \frac{ 4 }{ 9 }\]\[\Rightarrow \sin x = \sqrt{1-\frac{ 4 }{ 9 }}=\frac{ \sqrt{5} }{ 3 }\] \[\Rightarrow \tan x = \frac{ 3*\sqrt{5} }{ 4 }\] Is this correct or how does the condition 0 <a < pi/2 affect it?

OpenStudy (anonymous):

The interval 0 < x < pi/2 must simply tell the sign of the value of tan x right? Since tan > 0 for all values in 0 < x < pi / 2 the sign should be positive?

OpenStudy (ybarrap):

looks good, except you just have to square the 4/9 under the radical

OpenStudy (ybarrap):

Also, since your answer should be given in the range given, divide by pi (since tan x is periodic in period pi) to find how many periods back you need to subtract to get to that range.

OpenStudy (ybarrap):

To get to the range $$ 0\le x \le \pi/2 $$ Use this formula: \(\bf \Large x - \pi \times \lfloor\frac{x}{\pi}\rfloor\\ \text{ Where } \lfloor*\rfloor \text{ is the floor function}\\ \) I get \( \tan x = .82 \)

OpenStudy (ybarrap):

Another way to check if you answer is correct. Get a calculator and type tan(32.24) (use radians), then take the arctan of whatever you get. You should get .82.

OpenStudy (ybarrap):

Does this make sense?

OpenStudy (anonymous):

When calculating I get\[\tan x = \frac{\sqrt{65}}{4}, x = \arctan{\frac{\sqrt{65}}{4}} \approx 1.1102\] which is in the range [0,pi/2]